# Parameter estimation of range-migrating targets using OTFS signals from LEO satellites

### v2

@@ II-D OTFS demodulation
The radar receiver is implemented on a device than relies on its OTFS demodulator to obtain discrete samples of the received signal $y(t)$ in the delay-Doppler domain; such samples are then processed to estimate the target parameters as shown in Sec. V. The OTFS demodulator first transform $y(t)$ into the time-frequency domain via the Wigner Transform; in particular, the following measurements are obtained
@@
the cross-ambiguity function of $g_{\rm tx}(t)$ and $g_{\rm rx}(t)$; then, the following proposition establishes a connection between $\{Y_{\rm TF}[n,m]\}$ and $\{X_{\rm TF}[n,m]\}$ that generalizes the result in [8, Theorem 1] to the case where the variation of the target range is ignored only within each symbol interval.
@@
Proof.
Upon plugging (11) into (14), we have
which completes the proof. ∎
Finally, the OTFS demodulator transforms $Y_{\rm TF}[n,m]$ into the delay-Doppler domain via the Symplectic Finite Fourier Transform; in particular, we have
@@
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. In the Secs. III and IV we will establish an explicit relationship between $\{Y_{\rm DD}[n,m]\}$ and $\{X_{\rm DD}[n,m]\}$ for the cases where ideal and rectangular shaping filters are employed, respectively.
III Ideal shaping filters
Consider ideal shaping filters satisfying the bi-orthogonality condition [8, Sec. III]. Under this assumption, the ambiguity function satisfies $\gamma(\tau,\nu)=1$ for $(\tau,\nu)\in[-\tau_{\max},\tau_{\max}]\times[-\nu_{\max},+\nu_{\max}]$ and $\gamma(\tau,\nu)=0$ for $(\tau,\nu)\in[nT-\tau_{\max},nT+\tau_{\max}]\times[m\Delta-\nu_{\max},m\Delta+\nu_{\max}]$ for all pairs of integers $(m,n)\neq(0,0)$. Although practical waveforms can only approximate this condition to a certain extent, the use of ideal shaping filters offers valuable design insights, and the resulting system performance serves as a benchmark for comparison with that achievable using rectangular shaping filters.
In this case, (16) becomes
@@ III Ideal shaping filters
the periodic sinc function (also called Dirichlet function) of degree $Q$. Then, the following proposition establishes a connection between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ that generalizes the result in [8, Proposition 2] to the case where the target range can change every $N/B$ symbol intervals.
@@
Assume that ideal shaping filters are employed. Then, under the $B$-fold stop-and-go approximation on the range variation over the OFTS frame, we have
@@
Proof.
After exploiting (2) and (19), Eq. (18) becomes
Finally, notice that we have
$=NM\alpha\Phi[k-k^{\prime},l-l^{\prime}],$
where (25) follows from (7). ∎
III-A Some insights on Proposition 2
Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 0 of width $2/Q$ (between zero crossings) and Q2Q-2 sidelobes of width $1/Q$ (between zero crossings) in [0.5,0.5][-0.5,0.5]. The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [44]; following [8], for large $Q$, we can assume $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ is a design parameter that depends upon the desired approximation error. Accordingly, the following remarks are in order.
@@ III-A Some insights on Proposition 2
The sequence $\Phi[k,l]$ in (23) is a periodic function of $k$ and $l$ with period $N$ and $M$, respectively. This implies that (22) is a two-dimensional circular convolution, whereby
@@
$=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}\!X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big]\Phi[k^{\prime},l^{\prime}],$ (26b)
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. ∎
@@
For large $N/B$ and $M$, the sums in (26) only contain few dominant terms. To illustrate this point, denote by $R_{\rm r}=c/(M\Delta)$ and $R_{\rm rr}=\lambda/(NT)$ the range and range-rate resolutions [41], respectively, and write the range at time $bT_{B}$ and the range-rate as
@@
for $b=0,\ldots,B-1$, where $l_{b,{\rm int}}\in\{0,1,\ldots,M-1\}$, $k_{\rm int}\in\{0,1,\ldots,N-1\}$, and $l_{b,{\rm fra}},\,k_{\rm fra}\in(-0.5,0.5]$. Then, from (23) and (27) we have
@@
Finally, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Phi[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$: more on this in Remark 4. ∎
Remark 3.
Assume that $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=l_{{\rm fra}}$ for $b=0,\ldots,B-1$, which means that there is no range migration and that the target range is an integer multiple of the range resolution. Then, upon plugging (28) in (26b) and after straightforward manipulations, we obtain
consistently with [8, Proposition 2]. In addition, if $k_{\rm fra}=l_{\rm fra}=0$, then (29) becomes
Remark 4.
@@
Figure 2: $|\Phi[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ for Cases 1, 2, and 3, when $B=4$, $N=M=64$, and ideal shaping filters are employed.
The two-dimensional sequence $\alpha\Phi[k,l]$ represent the target response in the delay-Doppler domain. For a given $B$, $\Phi[k,l]$ is specified by the initial range and the range-rate of the target and satisfy $|\Phi[k,l]|\leq 1$; to get more insights on its structure, we discuss a toy example with $B=4$ and $N=M=64$. Fig. 2 shows $|\Phi[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ in different cases:
  1. 1.
    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0$, $k_{\rm int}=32$, and $l_{0,{\rm int}}=\cdots=l_{3,{\rm int}}=42$;
  2. 2.
    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$, $k_{\rm int}=32$, and $l_{0,{\rm int}}=\cdots=l_{3,{\rm int}}=42$;
  3. 3.
    $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{B-1,{\rm fra}}=0.5$, $k_{\rm int}=50$, $l_{0,{\rm int}}=42$, $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, and $l_{3,{\rm int}}=45$.
In Case 1, the range of the target is not changing in the OTFS frame; also, the range and range-rate of the target are multiple of $R_{\rm r}$ and $R_{\rm rr}$, respectively. In this case, $|\Phi[k,l]|$ only contains one non-zero entry, whose location depends upon the range and range-rate of the target, in keeping with the analytical model in (30). By comparing Cases 1 and 2, it is seen that a fractional range and a fractional range-rate cause a spread of $|\Phi[k,l]|$ along both domains, in keeping with (29). By comparing Cases 2 and 3, it seen that a variation of the target range every $N/B$ symbol intervals causes an additional spread of $|\Phi[k,l]|$ in both domains: this is the novel result established by Proposition 2. Finally, notice that most elements of $|\Phi[k,l]|$ are much smaller than one, in keeping with Remark 2; hence, the response of a high-speed target still has a sparse structure in the delay-Doppler domain, determined entirely by the initial range and range-rate. ∎
III-B Vectorized model
The samples $\{Y_{\rm DD}[k,l]\}$ can be organized into the vector $\bm{y}\in\mathbb{C}^{NM}$, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. In the following, we refer to $\bm{y}$ as the target echo. From (26a) we have
@@ III-B Vectorized model
where $\bm{\Phi}$ is an $NM\times NM$ matrix with $(\bm{\Phi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Phi\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{x}$ is an $NM$-dimensional vector with $(\bm{x})_{lN+k+1}=X_{\rm DD}[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. Alternatively, from (26b) we have
@@
where $\bm{X}$ is an $NM\times NM$ matrix with $(\bm{X})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=X_{\rm DD}\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{\phi}$ is an $NM$-dimensional vector with $(\bm{\phi})_{lN+k+1}=\Phi[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$.
In the following, we refer to
$\bm{e}(d,v,B)=\bm{\Phi}(d,v,B)\bm{x}=\bm{X}\bm{\phi}(d,v,B)$ (33)
as the normalized target echo, where we have made explicit the dependence upon the initial range $d$ of the target, the range-rate $v$ of the target, and the parameter $B$.
IV Rectangular shaping filters
Assume here $g_{\rm tx}(t)=g_{\rm rx}(t)=\sqrt{1/T}\mathrm{rect}(t/T)$. Then, $\gamma(\tau,\nu)=0$ for $|\tau|\geq T$; also, for large $M$, recall that [8]
@@ IV Rectangular shaping filters
To proceed, notice that (16) becomes here
@@
for $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, where the second and third terms on the right hand side are the inter carrier interference (ICI) and inter symbol interference (ISI) [8], respectively. For $n=0$ no ISI is present, as a guard interval is inserted between consecutive OTFS frames, and (16) becomes
@@
for $m=0,\ldots,M-1$. Then, the following proposition establishes a connection between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ that generalizes the result in [8, Theorem 2] to the case where the target range is not a multiple of $R_{\rm r}$ and can change every $N/B$ symbol intervals; the proof is in the Appendix.
@@
Assume rectangular shaping filters. Then, under the $B$-fold stop-and-go approximation on the range variation over the OFTS frame, we have
@@
The following remarks on Proposition 3 are now in order.
@@
The summations in (37) cannot be regarded as a two-dimensial convolution, since $\Psi_{k^{\prime},l^{\prime}}[k,l]$ depends upon $k^{\prime}$ and $l^{\prime}$: this is in sharp contrast with the result obtained for ideal shaping filters in Remark 1. Notice in passing that, for fixed $k^{\prime}$ and $l^{\prime}$, $\Psi_{k^{\prime},l^{\prime}}[k,l]$ is a periodic function of $k$ and $l$ with period $N$ and $M$, respectively, while, for fixed $k$ and $l$, it is a periodic function of $k^{\prime}$ with period $N$. ∎
@@
For large $N/B$ and $M$, the summations in (37) only contain few dominant dominant terms; indeed, from the properties of the Dirichlet function discussed in Sec. III-A, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$. This result is similar to that obtained for ideal shaping filters in Remark 2. ∎
Remark 7.
Proposition 3 subsumes [8, Theorem 2] as a special case. To verify this, use (27) to recast (37) as
@@
$\quad\times\Bigg[\sum_{l^{\prime}=0}^{M-1-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
@@
if $0\leq l\leq l_{{\rm int}}-1$, that is consistent with [8, Theorem 2]. ∎
@@
Following the discussion in Remark 4, the two-dimensional sequence $\alpha\Psi_{k^{\prime},l^{\prime}}[k,l]$ can be regarded as the target response in the delay-Doppler domain for an input symbol located on the delay-Doppler grid point indexed by $k^{\prime}$ and $l^{\prime}$. We underline that the target response depends now upon the location of the input symbol in the delay-Doppler plane, which is in sharp contrast to the result previously obtained for ideal shaping filters. For a given $B$, the sequence $\Psi_{k^{\prime},l^{\prime}}[k,l]$ in (38) is entirely specified by the initial range and the range-rate of the target and satisfy $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\leq 1$: this latter result is instead similar to that for ideal shaping filters. ∎
IV-A Vectorized model
As in Sec III-B, the samples $\{Y_{\rm DD}[k,l]\}$ can be organized into the vector $\bm{y}\in\mathbb{C}^{NM}$ representing the target echo, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. From (37), we have

### v3

@@ II-D OTFS demodulation
The radar receiver relies on an OTFS demodulator to obtain discrete samples of the received signal in the delay-Doppler domain. Such samples are then processed to estimate the target parameters as shown in Sec. IV.
First, the OTFS demodulator transforms $y(t)$ into the time-frequency domain via the Wigner Transform; in particular, the following measurements are obtained
@@
the cross-ambiguity function of $g_{\rm tx}(t)$ and $g_{\rm rx}(t)$; then, the following proposition establishes a connection between $\{Y_{\rm TF}[n,m]\}$ and $\{X_{\rm TF}[n,m]\}$ that generalizes the previous result in [6, Theorem 1]. The proof is reported in Appendix A.
@@
Next, the OTFS demodulator transforms $Y_{\rm TF}[n,m]$ into the delay-Doppler domain via the Symplectic Finite Fourier Transform; in particular, we have
@@
for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$.
III Delay–Doppler Input–Output Model
We establish here a relationship between $\{Y_{\rm DD}[k,l]\}$ and $\{X_{\rm DD}[k,l]\}$ for two representative shaping filters, namely ideal and rectangular filters. Ideal shaping filters lead to a simplified input–output relationship, and the resulting analysis provides useful insights into the impact of range migration and serves as a benchmark. In contrast, the more practical rectangular shaping filters yield a more involved relationship.
III-A Ideal shaping filters
Ideal shaping filters satisfies the bi-orthogonality condition [6, Sec. III]. In this case, the ambiguity function satisfies $\gamma(\tau,\nu)=1$ for $(\tau,\nu)\in[-\tau_{\max},\tau_{\max}]\times[-\nu_{\max},\nu_{\max}]$, and $\gamma(\tau,\nu)=0$ for $(\tau,\nu)\in[nT-\tau_{\max},nT+\tau_{\max}]\times[m\Delta-\nu_{\max},m\Delta+\nu_{\max}]$ for all integer pairs $(m,n)\neq(0,0)$. Consequently, (17) becomes
@@ III-A Ideal shaping filters
the periodic $\mathrm{sinc}$ function (also called Dirichlet function) [^7] Recall that $\mathcal{D}_{Q}(\nu)$ is a periodic function of $\nu$ with period one and is zero at multiple of $1/Q$, except for $\nu=0,\pm 1,\pm 2,\ldots$ where it is one; also, $|\mathcal{D}_{Q}(\nu)|$ has a main-lobe centered at 0 of width $2/Q$ (between zero crossings) and Q2Q-2 sidelobes of width $1/Q$ (between zero crossings) in [0.5,0.5][-0.5,0.5]. The first sidelobe is approximately $13$ dB down from the main-lobe, and the subsequent sidelobes fall off at about $6$ dB per octave [39]. of degree $Q$. Then, the following result extends [6, Proposition 2] to the case of target range variations every $N/B$ symbol intervals. The proof is reported in Appendix B.
@@
Assume that ideal shaping filters are employed. Under the blockwise constant-range approximation, we have
@@
The following remarks and example provide further insight into Proposition 2.
@@
$\Phi[k,l]$ is periodic in $k$ and $l$ with period $N$ and $M$, respectively. This implies that (23) is a two-dimensional circular convolution, which can be rewritten as
@@
$=\alpha\sum_{k^{\prime}=0}^{N-1}\sum_{l^{\prime}=0}^{M-1}X_{\rm DD}\Big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\Big]\Phi[k^{\prime},l^{\prime}].$ (25b)
Here, $\Phi[k,l]$ is the convolution kernel relating the transmitted and received delay-Doppler symbols. Accordingly, $\alpha\Phi[k,l]$ represents the target response in the delay-Doppler domain, and $\Phi[k,l]$ can be interpreted as the associated point-spread function, specified by the target initial range and range-rate.
@@
Proposition 2 subsumes [6, Proposition 2] as a special case. To illustrate this point, first use (1) to write the range at time $bT_{B}$ and the range-rate as
@@
where $l_{b,{\rm int}}\in\{0,1,\ldots,M-1\}$, $k_{\rm int}\in\{0,1,\ldots,N-1\}$, and $l_{b,{\rm fra}},\,k_{\rm fra}\in(-0.5,0.5]$. Next, assume no range migration, whereby $l_{b,{\rm int}}=l_{{\rm int}}$ and $l_{b,{\rm fra}}=l_{{\rm fra}}$ for $b=0,\ldots,B-1$. Upon plugging (26) in (25b), we obtain
consistently with [6, Proposition 2]. In addition, if $k_{\rm fra}=l_{\rm fra}=0$, then (27) becomes
Remark 3.
For large $N/B$ and $M$, the sum in (25) contains only a few dominant terms. Indeed, by substituting (26) into (24), we obtain
@@
For large $Q$, $|\mathcal{D}_{Q}(\nu)|\approx 0$ for $\nu\notin[-D/Q,D/Q]$, with $0<2D+1\ll Q$, where the non-negative integer $D$ controls the approximation error [6]. Accordingly, $|\Phi[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$.
@@
Figure 2: $|\Phi[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, when $B=4$, $N=M=64$, $l_{0,{\rm int}}=42$, $k_{\rm int}=32$, and ideal shaping filters. Case 1 (left): $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$ and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0$. Case 2 (center): $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$ and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0.5$. Case 3 (right): $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, $l_{3,{\rm int}}=45$, and $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=k_{\rm fra}=0.5$.
Example 1.
To provide physical insight into the effect of range migration on the target point-spread function in (24), we consider a toy example with $B=4$, $N=M=64$, $l_{0,{\rm int}}=42$, and $k_{\rm int}=32$. Fig. 2 reports $|\Phi[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ for different target kinematics. In Case 1, we assume no range migration (i.e., $l_{1,{\rm int}}=l_{2,{\rm int}}=l_{3,{\rm int}}=42$), integer range (i.e., $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0$), and integer range-rate (i.e., $k_{\rm fra}=0$). It is seen that $|\Phi[k,l]|$ contains a single non-zero entry at $(k,l)=(32,42)$, as predicted by (28). In Case 2, we assume no range migration, fractional range with $l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$, and fractional range-rate with $k_{\rm fra}=0.5$. By comparing Cases 1 and 2, it is observed that fractional range and range-rate induce a spreading of $|\Phi[k,l]|$ along both the delay and Doppler dimensions, as predicted by (27). In Case 3, we assume that the target range varies four times over the OTFS frame, with $l_{1,{\rm int}}=43$, $l_{2,{\rm int}}=44$, $l_{3,{\rm int}}=45$ and $k_{\rm fra}=l_{0,{\rm fra}}=\cdots=l_{3,{\rm fra}}=0.5$. By comparing Cases 2 and 3, range migration produces an additional spreading of $|\Phi[k,l]|$, which constitutes the main novelty captured by Proposition 2. In all cases, most entries of $|\Phi[k,l]|$ remain significantly smaller than one, in agreement with Remark 3. Hence, the response of a high-speed target preserves a sparse structure, albeit with an expanded support.
III-A1 Vectorized model
The samples $\{Y_{\rm DD}[k,l]\}$ can be arranged into the vector $\bm{y}\in\mathbb{C}^{NM}$, where $(\bm{y})_{lN+k+1}=Y_{\rm DD}[k,l]$, for $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$. From (25a), we obtain
@@ III-A1 Vectorized model
where $\bm{\Phi}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{\Phi})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=\Phi\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{x}\in\mathbb{C}^{NM}$ is defined by $(\bm{x})_{lN+k+1}=X_{\rm DD}[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. Alternatively, from (25b), we have
@@
where $\bm{X}\in\mathbb{C}^{NM\times NM}$ has entries $(\bm{X})_{lN+k+1,l^{\prime}N+k^{\prime}+1}=X_{\rm DD}\big[[k-k^{\prime}]_{N},[l-l^{\prime}]_{M}\big]$ and $\bm{\phi}\in\mathbb{C}^{NM}$ is defined by $(\bm{\phi})_{lN+k+1}=\Phi[k,l]$, for $k,k^{\prime}=0,\ldots,N-1$ and $l,l^{\prime}=0,\ldots,M-1$. Since the matrix $\bm{\Phi}$ and the vector $\bm{\phi}$ depend on the target initial-range and range-rate, we will denote them by $\bm{\Phi}(d,v)$ and $\bm{\phi}(d,v)$ when we want to make explicit this dependence.
III-B Rectangular shaping filters
Assume now $g_{\rm tx}(t)=g_{\rm rx}(t)=\sqrt{1/T}\,\mathrm{rect}(t/T)$. This implies that $\gamma(\tau,\nu)=0$ for $|\tau|\geq T$. Moreover, for large $M$, we have [6]
@@ III-B Rectangular shaping filters
Consequently, (17) simplifies as follows. For $n=1,\ldots,N-1$ and $m=0,\ldots,M-1$, we have
@@
where the second and third terms on the right-hand side correspond to ICI and ISI, respectively [6]. For $n=0$, instead, no ISI is present due to the guard interval between consecutive OTFS frames, and (17) reduces to
@@
for $m=0,\ldots,M-1$.
The following result extends [6, Theorem 2] to the case where the target range is not an integer multiple of $R_{\rm r}$ and may vary every $N/B$ symbol intervals. The proof is provided in Appendix C.
@@
Assume rectangular-shaped filters. Under the blockwise constant-range approximation, we have
@@
The following remarks and example provide further insight into Proposition 3.
@@
The right-hand side of (35) is not a two-dimensional convolution, since $\Psi_{k^{\prime},l^{\prime}}[k,l]$ depends on $k^{\prime}$ and $l^{\prime}$ due to ISI and ICI. This is in sharp contrast with the ideal shaping filter case. For fixed $k^{\prime}$ and $l^{\prime}$, $\Psi_{k^{\prime},l^{\prime}}[k,l]$ acts as a location-dependent kernel relating the input symbol $X_{\rm DD}[k^{\prime},l^{\prime}]$ to the output symbols $Y_{\rm DD}[k,l]$, and is periodic in $k$ and $l$ with period $N$ and $M$, respectively. Accordingly, $\alpha\Psi_{k^{\prime},l^{\prime}}[k,l]$ represents the target response in the delay-Doppler domain to an input symbol located at $(k^{\prime},l^{\prime})$, and $\Psi_{k^{\prime},l^{\prime}}[k,l]$ can be interpreted as a location-dependent point-spread function, specified by the target initial range and range-rate.
@@
By using (26), the result in (35) can be recast as
@@
$\quad\times\Bigg[\sum_{l^{\prime}=0}^{M-l_{b,{\rm int}}-\lfloor l_{b,{\rm fra}}\rfloor-1}X_{\rm DD}[k^{\prime},l^{\prime}]\Psi_{b,k^{\prime},l^{\prime}}[k-k^{\prime},l-l^{\prime}]$
@@
if $0\leq l\leq l_{{\rm int}}-1$, that is consistent with [6, Theorem 2].
@@
For large $N/B$ and $M$, the sum in (35) only contains few dominant terms; indeed, we have $|\Psi_{k^{\prime},l^{\prime}}[k,l]|\approx 0$ if $k-k_{\rm int}]_{N}\notin\{0,1,\ldots,DB\}\cup\{N-DB,\ldots,N-1\}$ or $l-l_{b,{\rm int}}]_{M}\notin\{0,1,\ldots,D\}\cup\{M-D,\ldots,M-1\}$, for $b=0,\ldots,B-1$. This result is similar to that obtained for ideal shaping filters in Remark 3.
Refer to caption    Refer to caption    Refer to caption
Refer to caption    Refer to caption    Refer to caption
Figure 3: $|\Psi_{k^{\prime},l^{\prime}}[k,l]|$ vs $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$, when $B=4$, $N=M=64$, and rectangular shaping filters are employed. Top and bottom rows are for $(k^{\prime},l^{\prime})=(0,0)$ and $(32,32)$, respectively. Cases 1 (left), 2 (center), and 3 (right) are defined as in Fig. 2.
Example 2.
Fig. 3 reports $|\Psi_{k^{\prime},l^{\prime}}[k,l]|$ versus $k=0,\ldots,N-1$ and $l=0,\ldots,M-1$ under the same parameter settings considered in Example 1, for the input locations $(l^{\prime},k^{\prime})=(0,0)$ and $(32,32)$. The resulting patterns show that the target responses exhibit similar overall support in the delay-Doppler domain, even though their shapes depend on the input location. For $(l^{\prime},k^{\prime})=(32,32)$, the Doppler profile is also affected by the term $\mathbbm{1}_{\{b=0\}}B/N$ in (38), which modifies the associated Dirichlet kernel due to the ISI induced by the rectangular shaping filters; this effect is not present for $(l^{\prime},k^{\prime})=(0,0)$.
III-B1 Vectorized model
The samples $\{Y_{\rm DD}[k,l]\}$ can again be arranged into the vector $\bm{y}\in\mathbb{C}^{NM}$. From (35), we obtain